2) A bag contains 10 beads. Six are black and 4 are white. ( i ) Three beads are selected from this bag at random, without replacement. What is the probability that precisely two of the beads are white? ( ii ) Three beads are selected from this bag, one at a time. After each bead is selected, its color is noted and it is replaced in the bag. What is the probability that precisely two of the beads are white?
Answer #1

i) P(two of the beads are white) = 4C2 * 6C1 / 10C3 = 6 * 6 / 120 = 0.3

ii) P(two of the beads are white) = 3C2 * (4/10)2 * (6/10) = 0.288